An aeroplane flies from origin to city located 175 kilometre 30 degree northof East next it flies 153 kilometre 20 degree centigrade west of north to city B.finally it flies 195 kilometre due West to city C find the total displacement
1
Expert's answer
2020-02-13T08:54:15-0500
"\bold{s_1}=(-175\cos{30},175\sin{30})=(-151.6,87.5)km"

"\bold{s_2}=(153\sin{20},153\cos{20})=(52.3,143.8)km"

"\bold{s_3}=(195,0)km"

"s_x=195+52.3-151.6=95.7\ km"

"s_y=87.5+143.8=231.3\ km"

The magnitude:


"s=\sqrt{95.7^2+231.3^2}=250\ km"

The direction:


"\theta=\arctan{\frac{95.7}{231.3}}=22.5\degree W\ of\ N"